# Polarized multiwavelength emission from pulsar wind – accretion disk interaction in a transitional millisecond pulsar

### v1

@@ Appendix C Timing Analysis of the IXPE Dataset
The top panel of Fig. 13 shows the IXPE background-subtracted double-peaked pulse profile. We calculate the fractional RMS variability amplitude () of the pulsed signal following the method by [121] and obtain = 8.1%0.8%. This value is consistent within the uncertainties with those derived in previous studies [5, 71, 98].
We obtained compatible results by analysing 0.5–6 keV NICER data in the high mode taken on 2024 June 1. The timing parameters had slightly larger uncertainties due to the lower number of photons recorded () compared to IXPE () and the uneven coverage.
@@
The spin frequency measured from IXPE data is compatible within the uncertainties with the value predicted from the quasi-coherent timing solution derived from optical data by [22] ( Hz), who found a spin down rate very close to the value observed when J1023 was detected as a radio pulsar before the state transition occurred in 2013. On the other hand, it differs by 16090 nHz from the value predicted by [71] ( Hz), who reported a spin-down rate about 20% faster (Fig. 11). Even though this difference is still compatible within a 3 c.l., our measurement tends to favor a scenario in which the rotational evolution of the pulsar in the subluminous X-ray state has remained largely unchanged compared to its behavior during the radio pulsar state.
Dataset and model summary | |
Pulsar name | J1023+0038 |
MJD range | 60459.5257—60475.6217 |
Data span (d) | 16.096 |
Solar system ephemeris | DE440 |
Time unit | TDB |
Parameters | |
POSEPOCH, Reference epoch for position () | 55000.00 |
RAJ, Right Ascension (J2000) | 10h23m47.687198s |
DECJ, Declination (J2000) | +00∘38′40.84551′′ |
F0, Spin frequency () | 592.42146705(9) |
PB, Orbital period () | 0.1980963155 |
A1, Projected semi-major axis of pulsar orbit (-) | 0.343356 |
TASC, Epoch of ascending node () | 60461.979580(14) |
Appendix D X-ray Polarimetric Properties
D.1 Model-independent Analysis
@@
We first perform a model-independent polarimetric analysis of the IXPE data using the formalism outlined by [76] implemented in the ixpeobssim package [11] under the pcube algorithm in the xpbin routine. We compute the normalized, background-subtracted Stokes parameters and , where , and are the Stokes parameters. From these, we derive the polarization degree and polarization angle (also referred to as the electric vector position angle) , measured counterclockwise from the North celestial pole toward the East. Hereafter, uncertainties are given at the 68.27% c.l. unless noted otherwise.
@@
In the 2–6 keV energy band, we obtain an MDP of 11% at a 99% c.l. and %. The polarization in the 2–3 keV energy range is consistent with zero within , while low-significance polarization is consistently detected across the 3–4, 4–5, and 5–6 keV energy ranges (Fig. 2). In the 3–6 keV range, we find an MDP of 15% and measure a polarization degree that exceeds this threshold, %, with . To assess the significance of this detection, we use the ratio of the polarization degree to its standard deviation, , as the primary indicator of statistical significance111111https://heasarc.gsfc.nasa.gov/docs/ixpe/analysis/IXPE_Stats-Advice.pdf. Using the distribution for two dofs, we calculate a confidence level (c.l.) of 99.4%.
@@
Applying the same analysis to high-mode data, we obtain an MDP of 12% in the 2–6 keV range, with % and . The energy-resolved analysis corroborates the results from the entire data set (Fig. 3). In the 3–6 keV range, we find an MDP of 16% and %, with .
@@
D.2 Spectro-polarimetric Analysis
@@
To enhance the significance of the polarimetric measurements, we conducted a spectro-polarimetric analysis of the data as outlined by [116], following the weighted analysis method prescribed by [38], where each photoelectron track recorded by the DUs is assigned a weight based on its ellipticity. We extracted the , and spectra for each DU using the xpbin tool’s PHA1 algorithms in ixpeobbsim, applying a constant energy binning of 200 eV. We assigned the response matrices 20240701-alpha075 available in the calibration database and computed modulation response functions and ancillary response files using ixpecalcarf.
@@
Previous observations have shown that the X-ray spectrum of J1023 during its high mode is well described by an absorbed power-law model (e.g., [18, 33, 24]). Hence, the adopted model includes a power-law component corrected for interstellar medium effects using the Tübingen-Boulder model [130] with photoelectric cross-sections by [124], convolved with an energy-independent polarization model. To account for potential discrepancies between the effective areas of the different DUs, we also included a cross-normalization constant, which was fixed at unity for DU1 to serve as the reference, and allowed to vary for DU2 and DU3. Thus, the final model employed within the XSpec spectral fitting package [6] is const TBabs powerlaw polconst. We fit this model simultaneously to the Stokes energy spectra of all DUs in the 2–6, 2–3 and 3–6 keV energy ranges using the forward-folding, maximum likelihood procedure implemented in XSpec.
@@
Observations over the past decade have shown that the X-ray spectral shape of J1023 has remained remarkably stable throughout this time span [18, 33, 24]. Hence, during the fitting process, we fixed the absorption column density () and power-law photon index () at the values measured in previous works: cm-2, . All other parameters were allowed to vary. The results are summarized in Table 1 and are consistent with those derived from the model-independent analysis. Figure 4 shows the confidence contours for the measurements of and obtained over distinct energy bands, drawn using the steppar command in XSpec.
@@
It is noteworthy that previous analyses of the distributions of mode durations extracted from observations carried out by XMM-Newton EPIC [5] show that 40% of low-mode episodes last less than 100 s. This implies that our method for selecting the emission modes inherently misses short low-mode episodes totaling 10% of the total duration of the IXPE observations. According to [10], the emission during the low mode is expected to be polarized to a lower extent compared to the high mode due to the much smaller contribution of the shock emission. Even in the most extreme case where the low-mode emission is fully unpolarized, the X-ray polarization degree during the high mode selected using our criteria would be enhanced by a factor of [40]. Here, and represent the background-subtracted count rates in the IXPE band in the low and high modes, respectively. This small increase, equivalent to about 0.2% in the polarization degree, is within the margin of the measurements uncertainties, meaning that as a matter of fact the intrinsic X-ray polarization degree of the high-mode emission remains unchanged within the limits of our observations.
D.3 Phase-Resolved Analyses
@@
We conducted both rotational and orbital phase-resolved analyses of the X-ray polarimetric properties. Using the photonphase tool in the PINT software package [86], we computed the rotational and orbital phases of each photon collected in the high mode in the 2–6 keV energy range by IXPE, based on the timing model derived from the IXPE data analysis. For each analysis, we split the data into 10 phase bins, and calculated the Stokes parameters for each bin using the pcube algorithm within ixpeobssim, following the same model-independent procedure described above and applying background subtraction.
@@
Figure 12 shows the Stokes parameters in distinct pulsar rotational phases (left) and binary orbital phases (right). The left panel of Figure 13 displays the Stokes parameters, and PAX,H as a function of the rotational phase. We do not measure statistically significant (anti-)correlations between the normalized rate and the Stokes parameters. Specifically, Spearman rank and Kendall tau correlation analyses between the normalized rate and yielded coefficients of -0.25 (p-value = 0.49) and -0.16 (p-value = 0.60), respectively. Similarly, correlation analyses between the normalized rate and resulted in coefficients of -0.47 (p-value = 0.17) and -0.33 (p-value = 0.22), respectively. Constant fits to the and values produced values of 8.5 and 11.8 for 9 dof, respectively. The right panel of Figure 13 shows the Stokes parameters, and PAX,H as a function of the orbital phase. Similarly, no significant variability was observed in and , with constant fits yielding values of 5.6 and 9.8 for 9 dof, respectively.
Appendix E Optical Properties
@@
All VLT observations resulted in significant detections (above 3), with a mean linear polarization level of (1). This value aligns with previous results obtained in the same band [8, 10], although it is slightly higher. The polarization angle is well-constrained, with a mean value of (1; measured after correcting for the polarization angle of the standard star). This value is consistent with that reported by [10], although our measurements are significantly more precise.
@@
In addition to the polarimetric analysis, it is also possible to extract and calibrate fluxes to build a light curve. If we assume the flux loss is negligible, the sum of the intensities measured in both the ordinary and extraordinary beams for each image gives the total flux from the target. Therefore, we summed the ordinary and extraordinary fluxes for each image of J1023, as well as for two isolated field stars with R2 magnitudes listed in the USNO B1 catalogue [91]. The resulting flux was then calibrated against the flux of the reference stars.
@@
Figure 1 shows the time evolution of the magnitude, Popt and PAopt. To quantify the intrinsic variability amplitude of the light curve, we computed the excess variance , where is the standard deviation of the magnitudes and is the mean square measurement error in magnitudes. We obtain ExcessVar 0.02 mag2, indicating moderate variability.
@@
During the low mode episode, we observe an increase in the optical emission by a factor of 1.3, which contrasts with the behavior seen in the X-ray and UV bands. The origin of this optical increase is unclear, but it may be related to the broadening or enhanced prominence of hydrogen H and/or three helium He I emission lines. These lines, possibly originating from regions within the accretion disk or from mass outflows, fall within the wavelength range of the filter of FORS2 (572.5–737.5 nm). A recent high-time-resolution optical spectroscopic study of J1023 using the OSIRIS spectrograph mounted on the GranTeCan reported significant variability in the equivalent width and full width at half maximum of these lines, such as changes by a factor of two over minute timescales for the H line, with no clear correlation with the flux of the optical continuum [89] (see also [110, 59] for further evidence of highly variable emission line profiles). In contrast, no prominent emission lines are expected within the wavelength range of the UVM2 filter of the Swift UVOT, as shown in the UV spectrum by [63]. We emphasize that if the optical emission lines originate from an outflow, this observation would lend support to scenarios suggesting stronger outflows during the low mode [122, 10], which also result in increased radio emission. However, since similar variability in the optical light curve is also observed in other instances during the X-ray high mode, it is not possible to draw a definitive conclusion.
@@
60465.9887 | 0.0131 0.0014 | -0.0009 0.0014 | 1.43 0.13 | 2.0 2.8 |
60465.9915 | 0.0127 0.0014 | 0.0009 0.0014 | 1.44 0.13 | -0.1 2.7 |
60465.9943 | 0.0101 0.0014 | 0.0005 0.0014 | 1.01 0.13 | 2.8 3.8 |
60465.9971 | 0.0074 0.0014 | -0.0037 0.0014 | 0.97 0.13 | -7.6 4.0 |
60465.9999 | 0.0129 0.0014 | -0.0012 0.0014 | 1.51 0.14 | -2.4 2.6 |
60466.0027 | 0.0090 0.0014 | -0.0044 0.0012 | 1.03 0.13 | -4.7 3.2 |
60466.0055 | 0.0131 0.0014 | -0.0069 0.0013 | 2.12 0.13 | -12.3 1.9 |
60466.0083 | 0.0113 0.0014 | -0.0041 0.0014 | 1.60 0.13 | -11.4 2.1 |
60466.0112 | 0.0101 0.0014 | -0.0018 0.0014 | 1.12 0.14 | -0.3 3.5 |
60466.0140 | 0.0129 0.0014 | 0.0021 0.0014 | 1.57 0.13 | 5.9 2.5 |
60466.0168 | 0.0138 0.0014 | -0.0028 0.0014 | 1.64 0.14 | -4.9 2.4 |
60466.0196 | 0.0152 0.0014 | -0.0030 0.0014 | 1.57 0.14 | 0.3 2.5 |
60466.0225 | 0.0108 0.0014 | 0.0023 0.0014 | 1.11 0.14 | -2.5 3.5 |
60466.0253 | 0.0106 0.0014 | -0.0037 0.0014 | 1.19 0.13 | -7.3 3.3 |
Popt exhibits moderate time variability, with a Median Absolute Deviation (MAD) of 0.23% and a normalized Excess Variance (NEV) of 0.05, indicating some intrinsic fluctuations. Its skewness is 0.58, suggesting a slight right-skewed distribution and indicating that deviations above the mean are more frequent. On the other hand, PAopt shows greater variability, with a MAD of 3.6∘, a significantly higher NEV of 12.7, and a skewness of -0.23, pointing to a slight left-skewed distribution where deviations below the mean are more common. The opposite trends in the skewness of Popt and PAopt are primarily due to their anticorrelated variability during the second peak of optical intensity (Figures 1 and 14). We measure and for the high mode; and for the low mode. Consistent with previous findings by [10], these measurements suggest a slight decrease in polarization degree when switching from the high to the low mode, while the polarization angle shows no significant variation within the uncertainties.
@@
Appendix F Polarization Properties in the Low Mode; Constraints on Polarized Radio Emission in the two Modes

### v2

@@ Appendix B Timing Analysis of the IXPE Dataset
The top panel of Fig. 11 shows the IXPE background-subtracted pulse profile. We calculate the fractional RMS variability amplitude () of the pulsed signal following the method by S. Vaughan et al. (2003) and obtain = 8.1%0.8%. This value is consistent within the uncertainties with those derived in previous studies (A. M. Archibald et al., 2015; A. Jaodand et al., 2016; A. Papitto et al., 2019).
@@
The spin frequency measured from IXPE data is compatible within the uncertainties with the value predicted from the quasi-coherent timing solution derived from optical data by A. Burtovoi et al. (2020) ( Hz), who found a spin down rate very close to the value observed when J1023 was detected as a radio pulsar before the state transition occurred in 2013. On the other hand, it differs by 16090 nHz from the value predicted by A. Jaodand et al. (2016) ( Hz), who reported a spin-down rate about 20% faster (Fig. 8). Even though this difference is still compatible within a 3 c.l., our measurement tends to favor a scenario in which the rotational evolution of the pulsar in the subluminous X-ray state has remained largely unchanged compared to its behavior during the radio pulsar state.
Appendix C X-ray Polarimetric Properties
C.1 Model-independent Analysis
@@
We first perform a model-independent polarimetric analysis of the IXPE data using the formalism outlined by F. Kislat et al. (2015) implemented in the ixpeobssim package (L. Baldini et al., 2022) under the pcube algorithm in the xpbin routine using the unweighted approach (A. Di Marco et al., 2022). We compute the normalized, background-subtracted Stokes parameters and , where , and are the Stokes parameters. From these, we derive the polarization degree and polarization angle (also referred to as the electric vector position angle) , measured counterclockwise from the North celestial pole toward the East. Hereafter, uncertainties are given at the 68.27% c.l. unless noted otherwise.
@@
In the 2–6 keV energy band, we obtain an MDP of 11% at a 99% c.l. and %. The polarization in the 2–3 keV energy range is consistent with zero within , while low-significance polarization is consistently detected across the 3–4, 4–5, and 5–6 keV energy ranges (Fig. 9). In the 3–6 keV range, we find an MDP of 15% and measure a polarization degree that exceeds this threshold, %, with . To assess the significance of this detection, we use the ratio of the polarization degree to its standard deviation, , as the primary indicator of statistical significance777https://heasarc.gsfc.nasa.gov/docs/ixpe/analysis/IXPE_Stats-Advice.pdf. Using the distribution for two dofs, we calculate a confidence level (c.l.) of 99.4%. Considering the 2–3 keV and 3–6 keV energy bins as independent data sets, we tested the combined significance of the polarization detection against the null hypothesis (i.e., zero polarization in every bin). This analysis yielded a polarization detection at a 96.1% c.l., corresponding to a probability for the null hypothesis.
@@
Applying the same analysis to high-mode data, we obtain an MDP of 12% in the 2–6 keV range, with % and . The energy-resolved analysis corroborates the results from the entire data set. In the 3–6 keV range, we find an MDP of 16% and %, with .
@@
C.2 Spectro-polarimetric Analysis
@@
To enhance the significance of the polarimetric measurements, we conducted a spectro-polarimetric analysis of the data as outlined by T. E. Strohmayer (2017) and adopted the weighted analysis method prescribed by A. Di Marco et al. (2022), where each photoelectron track recorded by the DUs is assigned a weight based on its ellipticity. We extracted the , and spectra for each DU using the xpbin tool’s PHA1 algorithms in ixpeobbsim, applying a constant energy binning of 200 eV. We assigned the response matrices 20240701-alpha075 available in the calibration database, which are appropriate for the time span covered by the observations, and computed modulation response functions and ancillary response files using ixpecalcarf.
@@
Previous observations have shown that the X-ray spectrum of J1023 during its high mode is well described by an absorbed power-law model (e.g., S. Bogdanov et al. 2015; F. Coti Zelati et al. 2018; S. Campana et al. 2019). Hence, the adopted model includes a power-law component corrected for interstellar medium effects using the Tübingen-Boulder model (J. Wilms et al., 2000) with photoelectric cross-sections by D. A. Verner et al. (1996), convolved with an energy-independent polarization model. To account for potential discrepancies between the effective areas of the different DUs, we also included a cross-normalization constant, which was fixed at unity for DU1 to serve as the reference, and allowed to vary for DU2 and DU3. Thus, the final model employed within the xspec spectral fitting package (K. A. Arnaud, 1996) is const TBabs powerlaw polconst. We fit this model simultaneously to the Stokes energy spectra of all DUs across the 2–6, 2–3 and 3–6 keV energy ranges using the standard forward-folding, maximum likelihood procedure implemented in xspec.
@@
Observations over the past decade have shown that the X-ray spectral shape of J1023 has remained remarkably stable throughout this time span (S. Bogdanov et al., 2015; F. Coti Zelati et al., 2018; S. Campana et al., 2019). Hence, during the fitting process, we fixed the absorption column density () and power-law photon index () at the values measured in previous works: cm-2, . All other parameters were allowed to vary. The results are summarized in Table 2 and are consistent with those derived from the model-independent analysis. Figure 2 shows the confidence contours for the measurements of and obtained over distinct energy bands, drawn using the steppar command in xspec.
@@
It is noteworthy that previous analyses of the distributions of mode durations extracted from observations carried out by XMM-Newton EPIC (A. M. Archibald et al., 2015) show that 40% of low-mode episodes last less than 100 s. This implies that our method for selecting the emission modes inherently misses short low-mode episodes totaling 10% of the total duration of the IXPE observations. According to M. C. Baglio et al. (2023), the emission during the low mode is expected to be polarized to a lower extent compared to the high mode due to the much smaller contribution of the emission from the boundary region. Even in the most extreme case where the low-mode emission is fully unpolarized, the X-ray polarization degree during the high mode selected using our criteria would be enhanced by a factor of (A. Di Marco et al., 2023). Here, and represent the background-subtracted count rates in the IXPE band in the low and high modes, respectively. This small increase, equivalent to about 0.2% in the polarization degree, is within the margin of the measurements uncertainties, meaning that as a matter of fact the intrinsic X-ray polarization degree of the high-mode emission remains unchanged within the limits of our observations.
C.3 Phase-Resolved Analyses
@@
We conducted both rotational and orbital phase-resolved analyses of the X-ray polarimetric properties. Using the photonphase tool in the PINT software package (J. Luo et al., 2021), we computed the rotational and orbital phases of each photon collected in the high mode in the 2–6 keV energy range by IXPE, based on the timing model derived from the IXPE data analysis. For each analysis, we split the data into 10 phase bins, and calculated the Stokes parameters for each bin using the pcube algorithm within ixpeobssim, following the same model-independent procedure described above and applying background subtraction.
@@
Figure 10 shows the Stokes parameters in distinct pulsar rotational phases. The left panel of Figure 11 displays the Stokes parameters, and PAX,H as a function of the rotational phase. We do not measure statistically significant (anti-)correlations between the normalized rate and the Stokes parameters. Specifically, Spearman rank and Kendall tau correlation analyses between the normalized rate and yielded coefficients of -0.25 (p-value = 0.49) and -0.16 (p-value = 0.60), respectively. Similarly, correlation analyses between the normalized rate and resulted in coefficients of -0.47 (p-value = 0.17) and -0.33 (p-value = 0.22), respectively. Constant fits to the and values produced values of 8.5 and 11.8 for 9 dof, respectively. The right panel of Figure 11 shows the Stokes parameters, and PAX,H as a function of the orbital phase. Similarly, no significant variability was observed in and , with constant fits yielding values of 5.6 and 9.8 for 9 dof, respectively.
Appendix D Optical Properties
@@
All VLT observations resulted in significant detections (above 3), with a mean linear polarization level of (1). This value aligns with previous results obtained in the same band (M. C. Baglio et al., 2016, 2023), although it is slightly higher. The polarization angle is well-constrained, with a mean value of (1; measured after correcting for the polarization angle of the standard star). This value is consistent with that reported by M. C. Baglio et al. (2023), although our measurements are significantly more precise.
@@
In addition to the polarimetric analysis, it is also possible to extract and calibrate fluxes to build a light curve. If we assume the flux loss is negligible, the sum of the intensities measured in both the ordinary and extraordinary beams for each image gives the total flux from the target. Therefore, we summed the ordinary and extraordinary fluxes for each image of J1023, as well as for two isolated field stars with R2 magnitudes listed in the USNO B1 catalogue (D. G. Monet et al., 2003). The resulting flux was then calibrated against the flux of the reference stars.
@@
Figure 1 shows the time evolution of the magnitude, Popt and PAopt. To quantify the intrinsic variability amplitude of the light curve, we computed the excess variance , where is the standard deviation of the magnitudes and is the mean square measurement error in magnitudes. We obtain ExcessVar 0.02 mag2. Converting this magnitude variability to flux variations, we derive an intrinsic fractional variability of 13%, indicating moderate variability.
@@
During the low mode episode, we observe an increase in the optical emission by a factor of 1.3, which contrasts with the behavior seen in the X-ray and UV bands. The origin of this optical increase is unclear, but it may be related to the broadening or enhanced prominence of hydrogen H and/or three helium He I emission lines. These lines, possibly originating from regions within the accretion disk or from mass outflows, fall within the wavelength range of the filter of FORS2 (572.5–737.5 nm). A recent high-time-resolution optical spectroscopic study of J1023 using the OSIRIS spectrograph mounted on the Gran Telescopio Canarias reported significant variability in the equivalent width and full width at half maximum of these lines, such as changes by a factor of two over minute timescales for the H line, with no clear correlation with the flux of the optical continuum (M. M. Messa et al., 2024) (see also T. Shahbaz et al. 2019; P. Hakala & J. J. E. Kajava 2018 for further evidence of highly variable emission line profiles). In contrast, no prominent emission lines are expected within the wavelength range of the UVM2 filter of the Swift UVOT, as shown in the UV spectrum by J. V. Hernandez Santisteban (2016). We emphasize that if the optical emission lines originate from an outflow, this observation would lend support to scenarios suggesting stronger outflows during the low mode (A. Veledina et al., 2019; M. C. Baglio et al., 2023), which also result in increased radio emission. However, since similar variability in the optical light curve is also observed in other instances during the X-ray high mode, it is not possible to draw a definitive conclusion.
Popt exhibits moderate time variability, with a Median Absolute Deviation (MAD) of 0.23% and a normalized Excess Variance (NEV) of 0.05, indicating some intrinsic fluctuations. Its skewness is 0.58, suggesting a slight right-skewed distribution and indicating that deviations above the mean are more frequent. On the other hand, PAopt shows greater variability, with a MAD of 3.6∘, a significantly higher NEV of 12.7, and a skewness of -0.23, pointing to a slight left-skewed distribution where deviations below the mean are more common. The opposite trends in the skewness of Popt and PAopt are primarily due to their anticorrelated variability during the second peak of optical intensity (Figures 1 and 12). We measure and for the high mode; and for the low mode. Consistent with previous findings by M. C. Baglio et al. (2023), these measurements suggest a slight decrease in polarization degree when switching from the high to the low mode, while the polarization angle shows no significant variation within the uncertainties.
@@
| 60465.9887 | 0.0131 0.0014 | -0.0009 0.0014 | 1.43 0.13 | 2.0 2.8 |
| 60465.9915 | 0.0127 0.0014 | 0.0009 0.0014 | 1.44 0.13 | -0.1 2.7 |
| 60465.9943 | 0.0101 0.0014 | 0.0005 0.0014 | 1.01 0.13 | 2.8 3.8 |
| 60465.9971 | 0.0074 0.0014 | -0.0037 0.0014 | 0.97 0.13 | -7.6 4.0 |
| 60465.9999 | 0.0129 0.0014 | -0.0012 0.0014 | 1.51 0.14 | -2.4 2.6 |
| 60466.0027 | 0.0090 0.0014 | -0.0044 0.0012 | 1.03 0.13 | -4.7 3.2 |
| 60466.0055 | 0.0131 0.0014 | -0.0069 0.0013 | 2.12 0.13 | -12.3 1.9 |
| 60466.0083 | 0.0113 0.0014 | -0.0041 0.0014 | 1.60 0.13 | -11.4 2.1 |
| 60466.0112 | 0.0101 0.0014 | -0.0018 0.0014 | 1.12 0.14 | -0.3 3.5 |
| 60466.0140 | 0.0129 0.0014 | 0.0021 0.0014 | 1.57 0.13 | 5.9 2.5 |
| 60466.0168 | 0.0138 0.0014 | -0.0028 0.0014 | 1.64 0.14 | -4.9 2.4 |
| 60466.0196 | 0.0152 0.0014 | -0.0030 0.0014 | 1.57 0.14 | 0.3 2.5 |
| 60466.0225 | 0.0108 0.0014 | 0.0023 0.0014 | 1.11 0.14 | -2.5 3.5 |
| 60466.0253 | 0.0106 0.0014 | -0.0037 0.0014 | 1.19 0.13 | -7.3 3.3 |
@@
Appendix E Polarization Properties in the Low Mode; Constraints on Polarized Radio Emission in the two Modes